OpenStudy (anonymous):

MEDAL AND FAN!!!! If the width of a rectangle is x + 2 and the length is twice as long, what is the area? A. 6x + 12 B. 2x² + 6x + 4 C. 2x² + 8x + 6 D. 2x² + 8x + 8

OpenStudy (anonymous):

c

OpenStudy (johnweldon1993):

The Area = Length x Width the width = x + 2 The length is twice that so 2(x + 2) So now that we have both length and width we have Area = 2(x + 2) * (x + 2) Can you take over from here?

OpenStudy (anonymous):

yea thnks you guys lol

OpenStudy (anonymous):

thanks for the metal and fan

OpenStudy (johnweldon1993):

Even though you really need to look that over again... @kawsam please refrain from giving answers, especially if you are unsure it they are correct...

OpenStudy (johnweldon1993):

I'll do this one out to show you... Area = 2(x + 2) * (x + 2) Area = (2x + 4) * (x + 2) 2x * x = 2x² 2x * 2 = 4x 4 * x = 4x 4 * 2 = 8 so we have 2x² + 4x + 4x + 8 2x² + 8x + 8 Again...please don't shoot out answers

OpenStudy (johnweldon1993):

@Rayvenn.14 see why I want you to do the questions out and not accept whatever anyone tells you...?

OpenStudy (anonymous):

yesss thanks you john

OpenStudy (johnweldon1993):

no problem

OpenStudy (anonymous):

thanks again